# Efficient generation of projective modules: a motivic view

### v1

@@ 2.4. Additional exact sequences involving $\pi_{j+1}\mathrm{St}_{n-j}({n})$
(21) $\mathrm{St}^{\mathrm{Sp}}_{2}({j})=\mathrm{Sp}_{j+2}/\mathrm{Sp}_{j}\simeq\mathrm{GL}_{j+2}/\mathrm{GL}_{j+1}\simeq\mathbb{A}^{j+2}\smallsetminus 0.$
@@ 3. Motivic obstruction theory and efficient generation
Let $k$ be an algebraically closed field and $X=\mathrm{Spec}\,R$ a smooth affine variety of dimension $d$. Murthy’s celebrated work implies that the $d$-th Segre class of a projective module $M$ of rank $r$ over $X$ is the only obstruction to generating $M$ by r+d1r+d-1 [15]. If we additionally assume that $k$ has characteristic zero, Murthy’s splitting conjecture in characteristic zero implies the following:
@@
Conversely, suppose that the $d$-th and (d1)(d-1)-st Segre class of $M$ vanishes and that $Q$ is a rank $d$ projective module such that $M\oplus Q$ is free (such an $Q$ exists by the Forster–Swan theorem). By hypothesis, the $d$-th and (d1)(d-1)-st Chern classes of $Q$ vanish. By Murthy’s conjecture in characteristic zero [2, Theorem 7.1.1], $Q\simeq Q^{\prime}\oplus R^{2}$ for some projective module $Q^{\prime}$ of rank d2d-2. We need a cancellation result to conclude that $M\oplus Q^{\prime}$ is free. For $r\geq 3$, this is Bass cancellation. For $r=2$, we appeal to Suslin’s celebrated work [17]. In the case $r=1$, we use Suslin’s Cancellation conjecture as resolved by Fasel [11, Theorem 2]. ∎
The main project of this section is to explore an obstruction-theoretic approach to reducing the number of generators of a projective module over smooth $k$-algebras when $k$ is not necessarily algebraically closed and not necessarily of characteristic zero. In Section 3.1 we recall the relevant obstruction theory. In Section 3.2, we review some classical vector bundle invariants, including Segre and Euler classes. In Section 3.3, we study efficient generation for line bundles. In Section 3.4 and Section 3.5, we identify certain key obstructions with Segre or Euler classes and give conditions for rank $r$ vector bundles on a smooth affine variety of Nisnevich cohomological dimension at most $d$ to be (r+d1)(r+d-1)- or (r+d2)(r+d-2)-generated.
@@ 3.1. Setting up obstruction theory for $\mathrm{Gr}_{r}({n})\to\mathrm{BGL}_{r}$
Let $X=\mathrm{Spec}\,R$ be a smooth affine variety of $\mathbb{A}^{1}$-cohomological dimension $d$ over a perfect field $k$ and let $n\geq r+2$. We consider the Moore-Postnikov factorization of the morphism $\mathrm{Gr}_{r}({n})\to\mathrm{BGL}_{r}$ representing the tautological rank $r$ bundle on $\mathrm{Gr}_{r}({n})$. We follow [5, Section 6.1] for the obstruction-theoretic set-up. By Proposition 2, this morphism fits into a fiber sequence
@@ 3.3. Efficient generation of line bundles and powers of the first Chern class
Let $k$ be an algebraically closed field. In [15, Corollary 3.16], Murthy shows that a rank $1$ module $L$ over a smooth affine $k$-algebra $R$ of dimension $d$ can be generated by $d$ elements if and only if the $d$-th power of the first Chern class of $L$ vanishes. In general, some hypothesis on $k$ will be necessary, but $k$ need not be algebraically closed and also need not satisfy other technical conditions listed in [15, Theorem 1.8].
@@ 3.4. The first nontrivial obstruction for rank at least $2$ and conditions for r+d1r+d-1 generation
Let $X=\mathrm{Spec}\,R$ be a smooth affine variety over a perfect field. Let $M$ be a projective module of rank $r\geq 2$. Suppose that $M$ can be generated by $n+1$ elements where $n\geq r+2$. Let $Q$ be a rank nr+1n-r+1 module such that $M\oplus Q\cong R^{n+1}$.
@@
where $\operatorname{pr}_{1}$ is projection onto the first factor and $h$ represents the surjection $M\oplus Q\to M$ from a free module of rank $n+1$ onto $M$. Our goal is to understand the potentially nontrivial obstruction to the existence of a dashed arrow, which is the obstruction to lifting to the (nr+1)(n-r+1)-st stage of the Moore–Postnikov tower.
We now relate the first obstructions in th Moore–Postnikov towers for the morphisms
@@
We relate $f$ from Equation 26 to maps we have already understood. We have a commutative diagram
(29) $\mathrm{St}_{r-1}({n})$$\mathrm{St}_{r}({n})$$\mathbb{A}^{n-r+1}\smallsetminus 0$$(\mathbb{A}^{n-r+1}\smallsetminus 0)\times\Omega(\mathbb{A}^{n+1}\smallsetminus 0),$$\scriptstyle{b_{n-r}^{r}}$$\scriptstyle{a_{n-r}^{r}}$$\scriptstyle{\,\,\operatorname{pr}_{1}}$$\scriptstyle{f}$
where the morphisms $a_{n-r}^{r}$ and $b_{n-r}^{r}$ are as in Definition 12. By Corollary 14, $f$ induces an isomorphism on $\pi_{n-r}$ for nrn-r even and is the quotient of $\mathrm{K}^{MW}_{n-r+1}$ by $\eta$ when nrn-r is odd.
Thus, for any parity of nrn-r, we find that $o_{n-r,n,r}(M)$ vanishes if $e_{n-r+1}(Q)=0$. If nrn-r is even, the converse is also true. In the case that nrn-r is odd, $o_{n-r,n,r}(M)$ is a unit multiple of $e_{n-r+1}(Q)$ modulo $\eta$, which is a unit multiple of the top Chern class of $Q$. Therefore $o_{n-r,n,r}(M)=0$ if and only if the top Segre class of $M$ vanishes. ∎
@@
In the statement of Proposition 29 with $n=r+1$, the Moore–Postnikov framework outlined in [5, Section 6.1] does not apply as stated to the study the problem of lifting along $\mathrm{Gr}_{r}({r+1})\to\mathrm{BGL}_{r}$, since the fiber $\mathrm{St}_{r}({r+1})$ is not simply connected. However, we can import the results of [16, Section 4] to the motivic setting. The action of $\pi_{1}\mathrm{BGL}_{r}\cong\mathbb{G}_{m}$ on $\pi_{1}\mathrm{St}_{r}({r+1})\cong\mathbb{G}_{m}$ arising from the fibration $\mathrm{St}_{r}({r+1})\to\mathrm{Gr}_{r}({r+1})\to\mathrm{BGL}_{r}$ is trivial, the first obstruction group is $\mathrm{CH}^{2}(X)$, and the first obstruction can again be identified with the Segre class. Given this, the remaining results in this section (Theorems 31, 32 and 33) apply when $d=2$, and those in the next section (Propositions 35, 36 and 37) apply when $d=3$. These modifications also prove the homotopy Forster–Swan theorem (Theorem 22) when $d=1$.
@@
If $k$ is algebraically closed and $R$ is a $k$-algebra of dimension $d$, then the previous corollary can be deduced from celebrated work of Murthy [15, Corollary 3.15]. On the other hand, note, for example, that Corollary 33 applies when $k$ is taken to be a finite field, since such fields have $2$-cohomological dimension equal to $1$.
@@ 3.5. Further reduction of the number of generators and secondary obstructions
By Proposition 29 with n=r+d2n=r+d-2 and $j=r$, the first obstruction to reducing the number of generators is $s_{r+d-1}(M)$ for $d$ odd and $e_{d+r-1}(Q)$ for $d$ even. The second obstruction is valued in $H^{d}(X,\pi_{d-1}\mathrm{St}_{r}({r+d-2})(\mathcal{L}))=0$, where $\mathcal{L}$ the determinant of $M$. ∎
@@
$g:=H^{d}(X,\pi_{d-1}g):H^{d}(X,\pi_{d-1}{\mathrm{St}_{r}({r+d-2})}(\mathcal{L}))\to H^{d}(X,\pi_{d-1}{\mathrm{St}_{r}({r+d-1})}(\mathcal{L}))$
on cohomology, where $\mathcal{L}=\det M$. We will show that $g$ is injective. Given this, the remaining obstruction to lifting along $A$ is precisely the first and only nontrivial obstruction to lifting along $B$, which by Corollary 33 is $s_{d}(M)$.
@@
where $b_{d-2}^{r}$ is as in Definition 12 with fiber $\mathbb{A}^{d-1}\smallsetminus 0$. We claim that the vertical morphism in Equation 32 induces an isomorphism after applying $H^{d}(X,\pi_{d-1}(-)(\mathcal{L})).$ Indeed, for $r\geq 3$, the vertical morphism is actually an isomorphism by Corollary 7. For $r=2$, the claim follows from Lemma 13 and [4, Proposition 5.2].
Now, consider the following commutative diagram:
(33) $H^{d}(X,\pi_{d-1}(\mathbb{A}^{d-1}\smallsetminus 0)(\mathcal{L}))$$H^{d}(X,\pi_{d-1}\mathrm{St}_{r}({r+d-2})(\mathcal{L}))$$H^{d}(X,\pi_{d-1}\mathrm{St}_{r-1}({r+d-2})(\mathcal{L}))$$H^{d}(X,\pi_{d-1}\mathrm{St}_{r}({r+d-1})(\mathcal{L})).$$\scriptstyle{g}\scriptstyle{\simeq}$
We claim the top row in Equation 33 is exact. Given this, if $H^{d}(X,\pi_{d-1}(\mathbb{A}^{d-1}\smallsetminus 0))=0$, then by [11, Lemma 2.2.3] the map $g$ is injective. To prove the claim, consider Proposition 15 with j=d2j=d-2, n=r+d2n=r+d-2. If $d$ is even, we see that the kernel of the map on $d$-th cohomology induced by $\pi_{d-1}b_{d-2}^{r}$ is $H^{d}(X,\operatorname{Im}(\pi_{d-1}a_{d-2}^{r})(\mathcal{L}))$ which is a quotient of $H^{d}(X,\pi_{d-1}(\mathbb{A}^{d-1}\smallsetminus 0)(\mathcal{L})).$ If $d$ is odd, consider again Proposition 15. Consider the sequence
(34) $H^{d-1}(X,\mathcal{I}^{d+1}(\mathcal{L}))$$H^{d}(X,\operatorname{Im}(\pi_{d-1}a_{d-2}^{r})(\mathcal{L}))$$H^{d}(X,\pi_{d-1}\mathrm{St}_{r}({r+d-2})(\mathcal{L}))$$H^{d}(X,2\mathrm{K}^{M}_{d}(\mathcal{L}))$$H^{d}(X,\mathrm{K}^{MW}_{d}(\mathcal{L}))$$H^{d}(X,\mathcal{I}^{d+1}(\mathcal{L}))$$\scriptstyle{\pi_{d-1}b_{d-2}^{r}\,\,\,\,\,\,}$
where $\mathcal{I}^{d+1}$ is the $d$-th power of the fundamental ideal in the Witt ring. Both the horizontal row and vertical column in Equation 34 are exact. Since $k$ is algebraically closed, $\mathcal{I}^{d+1}$ restricted to the small Nisnevich cite of $X$ is identically zero [4, Proposition 5.1] and the middle vertical morphisms is an isomorphism. ∎
@@ 4. A Forster–Swan theorem for symplectic modules
Let $X=\mathrm{Spec}\,R$ be a smooth affine over a perfect field $k$ and let $n\geq r+1$. Given an $R$ module $M$ of rank $2r$ equipped with a nondegenerate symplectic form $\omega\colon M\otimes M\to R$, one might seek an efficient generation result in the symplectic setting. The symplectic analogues of free modules are hyperbolic modules, i.e., direct sums of copies of the rank $2$ symplectic module $H$ given by $R^{2}$ with the form associated to the matrix
@@
Let $k$ be a perfect field of $2$-cohomological dimension at most $1$. Let $X=\mathrm{Spec}\,R$ be a smooth affine variety of dimension at most $d$ over $k$, and let $M$ be a symplectic module of rank $2r$ over $R$. If $d$ is even, then $M$ is a summand of a symplectic module of rank r+d2r+d-2 if and only if the Segre class $s_{d}(M)$ vanishes.
@@
Let $X=\mathrm{Spec}\,R$ be a smooth affine variety of odd dimension at most $d\geq 3$ over a perfect field $k$. Let $M$ be a rank $2r$ symplectic $R$-module.

### v2

@@ 2.4. Additional exact sequences involving $\pi_{j+1}\mathrm{St}_{n-j}({n})$
(21) $\mathrm{St}^{\mathrm{Sp}}_{2}({j+2})=\mathrm{Sp}_{j+2}/\mathrm{Sp}_{j}\simeq\mathrm{GL}_{j+2}/\mathrm{GL}_{j+1}\simeq\mathbb{A}^{j+2}\smallsetminus 0.$
@@ 3. Motivic obstruction theory and efficient generation
Let $k$ be an algebraically closed field and $X=\mathrm{Spec}\,R$ a smooth affine $k$-variety of dimension $d$. Murthy’s celebrated work implies that the $d$-th Segre class of a projective module $M$ of rank $r$ over $X$ is the only obstruction to generating $M$ by r+d1r+d-1 elements [24]. If we additionally assume that $k$ has characteristic zero, Murthy’s splitting conjecture in characteristic zero implies the following:
@@
Conversely, suppose that the $d$-th and (d1)(d-1)-st Segre class of $M$ vanishes and that $Q$ is a rank $d$ projective module such that $M\oplus Q$ is free (such an $Q$ exists by the Forster–Swan theorem). By hypothesis, the $d$-th and (d1)(d-1)-st Chern classes of $Q$ vanish. By Murthy’s conjecture in characteristic zero [2, Theorem 7.1.1], $Q\simeq Q^{\prime}\oplus R^{2}$ for some projective module $Q^{\prime}$ of rank d2d-2. We need a cancellation result to conclude that $M\oplus Q^{\prime}$ is free. For $r\geq 3$, this is Bass cancellation. For $r=2$, we appeal to Suslin’s celebrated work [27]. In the case $r=1$, we use Suslin’s Cancellation conjecture as resolved by Fasel [16, Theorem 2]. ∎
The main project of this section is to explore an obstruction-theoretic approach to reducing the number of generators of a projective module over smooth $k$-algebras when $k$ is not necessarily algebraically closed and not necessarily of characteristic zero. In Section 3.1 we explain the relevant obstruction theory. In Section 3.2, we review some classical vector bundle invariants, including Segre and Euler classes. In Section 3.3, we study efficient generation for line bundles. In Section 3.4 and Section 3.5, we identify certain key obstructions with Segre or Euler classes and give conditions for rank $r$ vector bundles on a smooth affine variety of Nisnevich cohomological dimension at most $d$ to be (r+d1)(r+d-1)- or (r+d2)(r+d-2)-generated.
@@ 3.1. Setting up obstruction theory for $\mathrm{Gr}_{r}({n})\to\mathrm{BGL}_{r}$
Let $X=\mathrm{Spec}\,R$ be a smooth affine variety of $\mathbb{A}^{1}$-cohomological dimension at most $d$ over a perfect field $k$ and let $n\geq r+2$. We consider the Moore-Postnikov factorization of the morphism $\mathrm{Gr}_{r}({n})\to\mathrm{BGL}_{r}$ representing the tautological rank $r$ bundle on $\mathrm{Gr}_{r}({n})$. We follow [5, Section 6.1] for the obstruction-theoretic set-up. By Proposition 2, this morphism fits into a fiber sequence
@@ 3.3. Efficient generation of line bundles and powers of the first Chern class
Let $k$ be an algebraically closed field. In [24, Corollary 3.16], Murthy shows that a rank $1$ module $L$ over a smooth affine $k$-algebra $R$ of dimension $d$ can be generated by $d$ elements if and only if the $d$-th power of the first Chern class of $L$ vanishes. In general, some hypothesis on $k$ will be necessary, but $k$ need not be algebraically closed and also need not satisfy other technical conditions listed in [24, Theorem 1.8].
@@
Remark 29.
The forward implication of Theorem 28 also follows from [17, Lemma 5.4], which in fact applies more generally to $X$ a smooth $k$-variety and $k$ an arbitrary field.
@@ 3.4. The first nontrivial obstruction for rank at least $2$ and conditions for r+d1r+d-1 generation
Let $X=\mathrm{Spec}\,R$ be a smooth affine variety over a perfect field. Let $M$ be a projective $R$-module of rank $r\geq 2$. Suppose that $M$ can be generated by $n+1$ elements where $n\geq r+2$. Let $Q$ be a rank nr+1n-r+1 module such that $M\oplus Q\cong R^{n+1}$.
@@
where $\operatorname{pr}_{1}$ is projection onto the first factor and $h$ represents the surjection $M\oplus Q\to M$ from a free module of rank $n+1$ onto $M$. Our goal is to understand the first potentially nontrivial obstruction to the existence of a dashed arrow, which is the obstruction to lifting to the (nr)(n-r)-th stage of the Moore–Postnikov tower.
We now relate the first obstructions in the Moore–Postnikov towers for the morphisms
@@
We now relate $f$ from Equation 26 to maps we have already understood. Note that $f\cong a_{r}^{n-r}\circ\operatorname{pr}_{1}$, with notation from Definition 12. So, by Corollary 14, $f$ induces an isomorphism on $\pi_{n-r}$ for nrn-r even and is the quotient of $\mathrm{K}^{MW}_{n-r+1}$ by $\eta$ when nrn-r is odd. Thus, for any parity of nrn-r, we find that $o_{n-r,n,r}(M)$ vanishes if $e_{n-r+1}(Q)=0$. If nrn-r is even, the converse is also true. In the case that nrn-r is odd, $o_{n-r,n,r}(M)$ is a unit multiple of $e_{n-r+1}(Q)$ modulo $\eta$, which is a unit multiple of the top Chern class of $Q$. Therefore $o_{n-r,n,r}(M)=0$ if and only if the top Segre class of $M$ vanishes. ∎
@@
In the statement of Proposition 30 with $n=r+1$, the Moore–Postnikov framework outlined in [5, Section 6.1] does not apply as stated to the study the problem of lifting along $\mathrm{Gr}_{r}({r+1})\to\mathrm{BGL}_{r}$, since the fiber $\mathrm{St}_{r}({r+1})$ is not simply connected. However, we can import the results of [25, Section 4] to the motivic setting. The action of $\pi_{1}\mathrm{BGL}_{r}\cong\mathbb{G}_{m}$ on $\pi_{1}\mathrm{St}_{r}({r+1})\cong\mathbb{G}_{m}$ arising from the fibration $\mathrm{St}_{r}({r+1})\to\mathrm{Gr}_{r}({r+1})\to\mathrm{BGL}_{r}$ is trivial, the first obstruction group is $\mathrm{CH}^{2}(X)$, and the first obstruction can again be identified with the Segre class. Given this, the remaining results in this section (Theorems 32, 33 and 34) apply when $d=2$, and those in the next section (Propositions 37, 38 and 39) apply when $d=3$. These modifications also prove the homotopy Forster–Swan theorem (Theorem 22) when $d=1$.
@@
If $k$ is algebraically closed and $R$ is a $k$-algebra of dimension $d$, then the previous corollary can be deduced from celebrated work of Murthy [24, Corollary 3.15]. On the other hand, note, for example, that Corollary 34 applies when $k$ is taken to be a finite field, since such fields have $2$-cohomological dimension equal to $1$.
In combination with results from [8], we also deduce efficient generation results for real algebraic vector bundles.
Theorem 36.
Let $X=\mathrm{Spec}\,R$ be a smooth affine algebra of dimension $d\geq 3$ over $\mathbb{R}$. Let $M$ be a rank $r\geq 2$ projective $R$-module over $R$. Then M can be generated by d+r1d+r-1 elements if and only if the top Segre class of M vanishes.
Proof.
The statement for $n$ even follows from Theorem 32. For $n$ odd, we appeal to [8, 2.1.1 and 2.1.2(3)]. ∎
@@ 3.5. Further reduction of the number of generators and secondary obstructions
By Proposition 30 with n=r+d2n=r+d-2 and $j=r$, the first obstruction to reducing the number of generators is $s_{d-1}(M)$ for $d$ odd and $e_{d-1}(Q)$ for $d$ even. The second obstruction is valued in $H^{d}(X,\pi_{d-1}\mathrm{St}_{r}({r+d-2})(\mathcal{L}))=0$, where $\mathcal{L}$ the determinant of $M$. ∎
@@
$g_{*}:=H^{d}(X,\pi_{d-1}g):H^{d}(X,\pi_{d-1}{\mathrm{St}_{r}({r+d-2})}(\mathcal{L}))\to H^{d}(X,\pi_{d-1}{\mathrm{St}_{r}({r+d-1})}(\mathcal{L}))$
on cohomology, where $\mathcal{L}=\det M$. We use analogous notation for other induced maps on cohomology. We will show that $g_{*}$ is injective. Given this, the remaining obstruction to lifting along $A$ is precisely the first and only nontrivial obstruction to lifting along $B$, which by Corollary 34 is $s_{d}(M)$.
@@
where $b_{d-2}^{r}$ is as in Definition 12 with fiber $\mathbb{A}^{d-1}\smallsetminus 0$. We claim that the vertical morphism in Equation 31 induces an isomorphism after applying $H^{d}(X,\pi_{d-1}(-)(\mathcal{L})).$ For $r\geq 3$, this is by Corollary 7. For $r=2$, the claim follows from Lemma 13 and [4, Proposition 5.2].
Now, consider the following commutative diagram, derived from Equation 31:
(32) $H^{d}(X,\pi_{d-1}(\mathbb{A}^{d-1}\smallsetminus 0)(\mathcal{L}))$$H^{d}(X,\pi_{d-1}\mathrm{St}_{r}({r+d-2})(\mathcal{L}))$$H^{d}(X,\pi_{d-1}\mathrm{St}_{r-1}({r+d-2})(\mathcal{L}))$$H^{d}(X,\pi_{d-1}\mathrm{St}_{r}({r+d-1})(\mathcal{L})).$$\scriptstyle{(b_{d-2}^{r})_{*}}$$\scriptstyle{g_{*}}\scriptstyle{\simeq}$
We claim the image of the horizontal arrow in Equation 32 contains the kernel of $g_{*}$. Given this, if $H^{d}(X,\pi_{d-1}(\mathbb{A}^{d-1}\smallsetminus 0))=0$, then by [16, Lemma 2.2.3] the map $g_{*}$ is injective. To prove the claim, consider Proposition 15 with j=d2j=d-2, n=r+d2n=r+d-2. If $d$ is even, we see that the kernel of the map on $d$-th cohomology induced by $\pi_{d-1}b_{d-2}^{r}$ is $H^{d}(X,\operatorname{Im}(\pi_{d-1}a_{d-2}^{r})(\mathcal{L}))$ which is a quotient of $H^{d}(X,\pi_{d-1}(\mathbb{A}^{d-1}\smallsetminus 0)(\mathcal{L})).$ If $d$ is odd, consider again Proposition 15. Consider the sequence
(33) $H^{d-1}(X,\mathcal{I}^{d}(\mathcal{L}))$$H^{d}(X,\operatorname{Im}(\pi_{d-1}a_{d-2}^{r})(\mathcal{L}))$$H^{d}(X,\pi_{d-1}\mathrm{St}_{r}({r+d-2})(\mathcal{L}))$$H^{d}(X,2\mathrm{K}^{M}_{d}(\mathcal{L}))$$H^{d}(X,\mathrm{K}^{MW}_{d}(\mathcal{L}))$$H^{d}(X,\mathcal{I}^{d}(\mathcal{L}))$$\scriptstyle{(b_{d-2}^{r})_{*}\,\,\,\,\,\,\,}$
where $\mathcal{I}^{d}$ is the $d$-th power of the fundamental ideal in the Witt ring. Both the horizontal row and vertical column in Equation 33 are exact. Since $k$ is algebraically closed and $d\geq 4$, [4, Proposition 5.2] implies that the middle vertical morphism is an isomorphism. ∎
@@ 4. A Forster–Swan theorem for symplectic modules
Let $X=\mathrm{Spec}\,R$ be a smooth affine variety over a perfect field $k$ and let $n\geq r+1$. Given an $R$ module $M$ of rank $2r$ equipped with a nondegenerate symplectic form $\omega\colon M\otimes M\to R$, one might seek an efficient generation result in the symplectic setting. The symplectic analogues of free modules are hyperbolic modules, i.e., direct sums of copies of the rank $2$ symplectic module $H$ given by $R^{2}$ with the form associated to the matrix
@@
Let $k$ be a perfect field of $2$-cohomological dimension at most $1$. Let $X=\mathrm{Spec}\,R$ be a smooth affine variety of dimension at most $d$ over $k$, and let $M$ be a symplectic module of rank $2r$ over $R$. If $d$ is even, then $M$ is a summand of a symplectic module of rank 2r+d22r+d-2 if and only if the Segre class $s_{d}(M)$ vanishes.
@@
Let $X=\mathrm{Spec}\,R$ be a smooth affine variety of odd dimension at most $d\geq 3$ over a perfect field $k$. Let $M$ be a rank $2r\geq 4$ symplectic $R$-module.